Opened 12 years ago

Last modified 6 years ago

#342 closed flight processing

GB08/06, flight day 167/2010, Bleaklow — at Version 30

Reported by: ella Owned by: jb362
Priority: alpha 4 low Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: iopa@… Other processors:

Description (last modified by iopa)

Location of unpacked data: /users/rsg/arsf/workspace/GB08_06-2010_167_Bleaklow

Scientific purpose: Study of peatlands in the uplands of the Peak District National Park.

Scientific objectives: To evaluate the ability of SPECIM images to distinguish between moorland restoration treatments and provide new
knowledge on thresholds for estimating habitat condition for sparse vegetation growth on carbon-rich soils, building on work on hyperspectral remote sensing of bare peat.

Data arrived from ARSF via network transfer on 28th June 2010.

PI: Ms Julia McMorrow

Sensors:

  • Eagle (18/08/10)
  • Hawk
  • Leica LIDAR (18/08/10)
  • RCD Digital photography (24/08/10)

Change History (28)

comment:1 Changed 12 years ago by ella

  • Component changed from ARSF to Processing: general

comment:2 Changed 12 years ago by benj

  • Summary changed from BG08/06, flight day 167/2010, Bleaklow to GB08/06, flight day 167/2010, Bleaklow

comment:3 Changed 12 years ago by benj

  • Description modified (diff)

comment:4 Changed 12 years ago by ella

This project has not yet been fast copied into the arsf repoistry.

comment:5 Changed 12 years ago by ella

Applanix navigation processing done.

comment:6 Changed 12 years ago by ella

IPAS navigation processing is done.

comment:7 Changed 12 years ago by jb362

  • Owner changed from mggr to jb362
  • Status changed from new to assigned

Starting hyperspectral processing.

comment:8 Changed 12 years ago by jb362

North end of eagle flightline 8 will not line up with other flightlines.The 0.03 offset seems to be the best fit for the rest of the flightline.

comment:9 Changed 12 years ago by jb362

Playing around with the gpt offsets doesn't appear to sort out the above problem

comment:10 Changed 12 years ago by jb362

I think I may just cut off the top of flight line 8 off since changing the sct or gpt offests doesn't have any effect

comment:11 Changed 12 years ago by benj

No, you shouldn't ever cut data off the flightline. If it can't be rectified completely then put a comment to that effect in the readme prior to delivery, but the data itself is still fine even if not quite in the right place.

comment:12 Changed 12 years ago by jb362

OK will do.

SCT Offsets:

e1: 0.03
e2: 0.01
e3: 0.04
e4: 0.01
e5: 0.00
e6: 0.01
e7: -0.01
e8: 0.03
e9: 0.03
e10: 0.08
e11: 0.05
e12: -0.01
e13: 0.05
e14: 0.02
e15: 0.03
e16: 0.85
e17: 0.01
e18: -0.01
e19: 0.03
e20: 0.04
e21: 0.02
e22: 0.02
e23: 0.05
e24: 0.04
e25: -0.01
e26: 0.03

Sync error on flightline 16 accounts for very large offset

comment:13 Changed 12 years ago by mark1

OS vectors have arrived and have been unpacked to ~arsf/vectors/GB08_06-second

comment:14 Changed 12 years ago by jb362

Flight line e23 split into 4 parts:

e24a:0.05
e24b:0.1
e24c:0.1
e24d:0.06

Flight lines e11, e21, e22 had end of flight line cut off as these sections were out the target area and had large bends resulting in turn errors.

comment:15 Changed 12 years ago by jb362

Typo!

Flight line e23 split into 4 parts:

e23a:0.05
e23b:0.1
e23c:0.1
e23d:0.06

comment:16 Changed 12 years ago by iopa

  • Cc iopa@… added

Taking up LIDAR processing.

comment:17 Changed 12 years ago by jb362

Hawk SCT offsets:

h1: 0.02
h2: 0.01
h3: 0.06
h4: 0.00
h5: 0.03
h6: 1.70
h7: 0.01
h8: 0.02
h9: 0.00
h10: 0.01
h11: 0.02
h12: -0.01
h13: 0.02
h14: 0.03
h15: 0.02
h16: 1.35
h17: 0.00
h18: 0.03
h19: 0.01
h20: 0.03
h21: 0.02
h22: 0.03
h23: 0.02
h24: 0.05
h25: 0.06
h26: 0.01

Sync errors on flight lines 16 & 6 account for very large offsets

comment:18 Changed 12 years ago by iopa

Missing .sup file from IPAS output, will re-process IPAS before proceeding to LIDAR.

comment:19 Changed 12 years ago by jb362

Hyperspectral processing complete and ready for delivery check

comment:20 Changed 12 years ago by anch

Delivery checking the hyperspectral data for this project now.

comment:21 Changed 12 years ago by jb362

Delivery check fixes complete

comment:22 Changed 12 years ago by jb362

Delivery checking LiDAR

comment:23 Changed 12 years ago by jb362

Check complete. Change "Ordance survey" to "Ordnance survey" in data quality remarks.

comment:24 Changed 12 years ago by iopa

Dispatched USB disk containing Eagle and LiDAR on the 19/08/2010.
PI and collaborators have been emailed.

Note: Collaborator 2 will be unavailable and would like any contact regarding remote sensing to be diverted to Stuart Masheder.

Hawk data will be delivered individually when calibration issue has been resolved.

comment:26 Changed 12 years ago by iopa

  • Description modified (diff)

Processing camera.

comment:27 Changed 12 years ago by iopa

Following a request by the PI, first collaborator Aletta Bonn no longer works on the project.
In her stead, Jonathan Walker will act as first collaborator .

Also by request of the PI, Elizabeth Lowe should be CC'ed during any contact.(emails can be found in Internal Contact details page)

comment:29 Changed 12 years ago by jb362

Camera delivery check complete- everything looks good.

comment:30 Changed 12 years ago by iopa

  • Description modified (diff)
Note: See TracTickets for help on using tickets.